### v1

# Untestability of Average Slutsky Symmetry

Abstract
Slutsky symmetry and negative semidefiniteness are necessary and sufficient conditions for the rationality of demand functions. While the empirical implications of Slutsky negative semidefiniteness in repeated cross-sectional demand data are well understood, the empirical content of Slutsky symmetry remains largely unexplored. This paper takes an important first step toward addressing this gap. We demonstrate that the average Slutsky matrix is not identified and that its identified set always contains a symmetric matrix. A key implication of our findings is that the symmetry of the average Slutsky matrix is untestable, and consequently, individual Slutsky symmetry cannot be tested using the average Slutsky matrix.
Keywordsβ Slutsky symmetry, rationality, cross-sectional demand, continuity equation
1 Introduction
Rationality is a central concept in economic theory, serving as a fundamental assumption in the analysis of consumer behavior. It assumes that consumers make decisions consistent with maximizing utility, a principle that supports many economic models and empirical research. Testing this assumption is crucial for validating theoretical models and understanding consumer decision-making.
When individual demand functions are available, the Hurwicz-Uzawa theorem ([hurwicz1971integrability]) provides a complete characterization of rationality. To state the theorem, let be a -dimensional price vector and be an income, both of which are relative to the price of the numeraire. Let be the -dimensional vector of quantities demanded when the price is and the income is We say that a demand function is rational if it maximizes a utility function subject to the budget constraint. According to the Hurwicz-Uzawa theorem, a demand function is rational if and only if its Slutsky matrix, which is defined as a matrix
is both symmetric and negative semidefinite.
In many real-world applications, individual demand functions are not observed and only cross-sectional data is available. To investigate the population rationality, researchers have explored the rationalizability of the average demand function conditional on observable characteristics, such as price and income, treating it as if it is generated by an individual representing the population. See, for example, [lewbel1995consistent], [lewbel2001demand], [haag2009testing], and [hoderlein2011many]. However, this approach involves some disadvantages. For instance, as Theorem 1 of [lewbel2001demand] demonstrates, the rationality of the average demand function is unrelated to the rationality of individual demand functions: it is possible for all individuals in a population to be rational while the average demand is not, and vice versa. Moreover, it is often reported that cross-sectional mean regressions fail to explain the variation of demand adequately because of the unobserved preference heterogeneity. ([hoderlein2011many], [hausman2016individual].) This fact indicates the need to investigate all the information about the heterogeneity contained in the data.
Testing rationality for cross-sectional demand data is a harder problem than that for individual demands because individual Slutsky matrices are not observed. This difficulty raises a fundamental question: can we infer the rationality of a population from such data? More specifically, is it possible to construct statistical tests with nontrivial power to detect consumer irrationality using only distributional data?
Recent research has made progress in addressing this question. [hausman2016individual] provide a necessary and sufficient condition for cross-sectional demand distributions to be rationalizable when only two goods are present. For cases involving more than two goods, [dette2016testing] provide the empirical content of the negative semidefiniteness of the Slutsky matrix. Specifically, they show that the quadratic form of the average Slutsky matrix conditional on price and income, where is the random demand function representing the population, is identified. If an estimate of the quadratic form is significantly positive, then it means that the observed data is likely to be inconsistent with a population having an negative semidefinite Slutsky matrixβlet alone rationality. These studies demonstrate the feasibility of testing rationality by focusing on the negative semidefiniteness of the Slutsky matrix. However, these methods neglect the symmetry condition, leading to statistical tests that are overly conservative and fail to fully leverage the implications of rationality. The symmetry condition can be more critical to the rationality than the negative semidefiniteness, given that the former is highly sensitive to small perturbations while the latter is not. If Slutsky symmetry is testable, one should be able to construct a much more powerful test for rationality by testing both conditions.
Despite its theoretical importance, the empirical content of the Slutsky symmetry condition remains poorly understood. This paper provides a first step toward elucidating the implications of individual Slutsky symmetry for cross-sectional demand distributions. We demonstrate that the average Slutsky matrix is not identified unlike its quadratic form, which is identified as shown by [dette2016testing]. Moreover, its identified set always contains a symmetric matrix. To prove these results, we construct a stochastic demand system that is observationally equivalent to the true demand system while ensuring that the average Slutsky matrix is symmetric.
An immediate consequence of our result is the fundamental untestability of individual Slutsky symmetry via the average Slutsky matrix. This fact indicates that the Slutsky symmetric condition is totally distinct to the negative semidefiniteness condition, which is testable based on the quadratic form of the average Slutsky matrix ([dette2016testing]). Note also that it does not immediately rule out the possibility of testing Slutsky symmetry from cross-sectional distributional data. It remains an open question whether individual symmetry has observable implications through nonlinear statistics, a topic that is not explored in this paper.
The key theoretical contribution of this paper is the introduction of the continuity equation as a tool to construct a random demand function that satisfies given marginal distributions and adheres to conditions on the Slutsky matrix. The continuity equation originates from fluid dynamics, where it describes the time evolution of fluid density. It is expressed as:
where is the divergence operator defined as for a vector-valued function is the fluid density, and is the flow velocity vector field at time The first term of the LHS represents the change in fluid volume within the system, while the second term accounts for the net flow difference (inflow minus outflow). This equation describes situations where all changes in fluid volume are due to particle transport, with no external sources or sinks present. In the context of demand analysis, heterogeneous individuals are conceptualized as a fluid, with price and income interpreted as βmultidimensional time.β The continuity equation then captures how demand distributions evolve as price and income change. The velocity field corresponds to the derivatives of demand with respect to price and income, which are components of the Slutsky matrix. Our approach to constructing a desired random demand function involves finding a βvelocityβ field that obeys the continuity equation while satisfying the rationality condition.
Related literature. [hurwicz1971integrability] investigate the integrability of individual demand functions and give a necessary and sufficient condition for rationality based on the Slutsky matrix. [lewbel2001demand] considers a population that is heterogeneous in preference and provides conditions for the average demand function to be rational. [haag2009testing] and [hoderlein2011many] among others consider the estimation and inference on the average demand function under the rationality shape restriction.
For cross-sectional demand distributions, [hausman2016individual] give a necessary and sufficient condition for observed datasets to be consistent with a rational demand system for the cases when there are only two goods. When more than two goods are present, the characterization of rationalizability is largely open. [dette2016testing] provide a necessary condition that the data needs to satisfy for it to be rationalizable focusing on the negative semidefiniteness of the Slutsky matrix and propose a statistical testing for rationality. [maes2024beyond] construct a testing procedure of rationality based on higher order moments of demand distributions. They also observe that the average Slutsky matrix is not identified from those moments.
This paper is also related to the literature on random utility models. [mcfadden1990stochastic] and [mcfadden2005revealed] show that the Axiom of Revealed Stochastic Preference characterizes rationalizability of stochastic choice functions defined on a finite number of choice sets. In a similar setup, [kitamura2018nonparametric] constructed a statistical test for the axiom.
In the context of discrete choice, [bhattacharya2025integrability] recently gives the complete characterization of the rationalizability of demand distributions.
2 Setup and Main Result
For we consider an economy with goods. Relative to the first good, their prices are encoded into a price vector where Let be the relative income. For given price and income a consumer demands Notice that we assume the homogeneity of demand functions at this point. We also assume Warlasβ law, i.e., the demand for the numeraire is and consequently, is assumed.
We assume
for and Let We restrict ourselves to demand functions that are in the space
In the real world, consumerβs preference heterogeneity is present. In this sense, consumerβs demand is stochastic from the perspective of an econometrician. Let be a random individual demand function drawn from a probability distribution on The econometrician is assumed to observe cross-sectional demand distributions, that is, (s)he observes the distribution of for each but no joint distribution of demands across different price-income levels is available. Regarding as a stochastic process indexed by we often call the (one-dimensional) marginal distribution of at Note that although is a population object, we assume that it is available since we are interested in identification. Assume that the interior of the support of denoted by is not empty.
Example 1.
For the three-good case (), consider a random Cobb-Douglas demand where is a random vector such that and Then is a -valued random element. The demand distribution conditional on is a distribution supported on a subset of the triangle generated by and
Recall that an individual demand function is said rational if it is induced by utility maximization. The Hurwicz-Uzawa theorem states that is rational if and only if its Slutsky matrix
is symmetric and negative semidefinite for each
As we discussed in the previous section, [dette2016testing] propose a method to test individual rationality based on the negative semidefiniteness. Their testing procedure is roughly as follows. In their Theorem 1, they show that the quadratic form of the average Slutsky matrix is identified, that is, is identified for each and If the population consists of rational individuals, then is negative semidefinite almost surely, and hence, should hold. One can reject the null hypothesis that the population is rational if an estimate of is significantly positive for some and 333To be more precise, [dette2016testing] show a stronger result that the quadratic form of the average Slutsky matrix conditional on the value of Thus, their test is more powerful than what is described here.
Although this method has nontrivial power, it is likely to be overly conservative since it completely neglects the other critical component of rationality, Slutsky symmetry. We shall investigate its testability based on the average Slutsky matrix. This question is important since the symmetry property is not robust to small perturbations while the negative semidefiniteness is. If Slutsky symmetry is testable, one should be able to construct a much more powerful test for rationality by testing both conditions.
Unfortunately, the following theorem shows that Slutsky symmetry is not testable based on the average Slutsky matrix. The proof is illustrated in the next section.
Theorem 2.1.
Suppose satisfies Assumption B.1. There exists a -valued random demand function such that and is symmetric for all
Theorem 2.1 particularly implies that the average Slutsky matrix is not identified in our setup. The difficulty arises from the second term of (1). The expectation of the second term is written as
but it is unclear how to identify because it involves the joint distribution of demand at different income levels, and it is indeed not identifiable as we see in the proof of Theorem 2.1.
Theorem 2.1 implies that no matter what cross-sectional demand distributions satisfying the regularity condition we observe, there always exists a stochastic demand system such that it is observationally equivalent to the true demand system and its average Slutsky matrix is symmetric. Hence, there is no way to infer whether is symmetric or not from
Theorem 2.1 does not immediately rule out the possibility of testing Slutsky symmetry from cross-sectional distributional data. It can be still possible that individual symmetry has observable implications through nonlinear statistics, but we leave this for future work.
Remark 2.1.
In addition to the marginal distribution of let us assume that the joint distribution of is available for all This setup corresponds to the situation where we can observe each individualβs choice twice. In this case, the average Slutsky matrix is identified because the RHS of (2) is identified. Hence, it is possible to test the Slutsky symmetry by testing whether is symmetric or not.
Remark 2.2.
Proposition 2 of [maes2024beyond] asserts that the average Slutsky matrix is not βautomaticallyβ identified. They explain that the quantity (2) is not identified in the same way that the other term is identified through the equation Their argument is not complete, as it merely rules out a particular strategy for identifying the average Slutsky matrix without addressing the possibility of alternative identification approaches. Consequently, [maes2024beyond] do not establish whether there exists an observationally equivalent demand system whose average Slutsky matrix is symmetric.
3 Construction of Stochastic Demand Systems
In this section, we illustrate how a stochastic demand system with the properties of Theorem 2.1 is constructed for the three-good case (). General cases are discussed in the appendix. Remember that for given we need to construct a -valued random element such that and is symmetric. We first focus on the first condition, the marginal compliance. Let
Lemma 3.1.
Let be such that Assumption B.1. Then, there exists a measurable function such that
1.
it is continuously differentiable in each argument,
2.
is a homeomorphism for each
3.
for each and
4.
for each
In particular, the -valued random function where complies the marginal distributions,
In the random demand function constructed in Lemma 3.1, consumersβ preference heterogeneity, or βtype,β is encoded in By the third property of consumerβs type is understood as the demand at Observe that the demand system constructed in Lemma 3.1 is degenerated in the sense that consumers are completely characterized by the demand at More specifically, if an individual demands at (s)he demands at for sure.
Although in Lemma 3.1 satisfies the marginal compliance, it does not satisfy the average Slutsky symmetry condition in general. Our strategy is to construct another random demand function satisfying both conditions by modifying The following lemma gives a condition equivalent to the symmetry.
Lemma 3.2.
Let be such that Assumption B.1. Suppose that a -valued random function satisfies for all Then, its average Slutsky matrix is symmetric if and only if
holds.
We shall modify so that it satisfies equation (3). Fix Let be the income derivative of the demand function of consumer
where Notice that this is well-defined since is homeomorphic. The family of vector fields, combined with an initial condition, pins down the demand function, as is the unique solution to the ODE
under the global Lipschitz condition on by the Cauchy-Lipschitz theorem. Since is not equal to the RHS of (3) in general, we shall rectify with an auxiliary vector field constructed in the following lemma.
Lemma 3.3.
Fix Let be such that is smooth. There exists a vector field that is Lipschitz uniformly in and satisfies
For
solve the partial differential equation (PDE) (4) and let Consider the following ODE
By the Cauchy-Lipchitz theorem, this ODE admits a unique solution for each Define
Lemma 3.4.
The random function constructed in this lemma satisfies the conditions of Theorem 2.1.
In a similar idea, we can explicitly write down the identified set of the average Slutsky matrix.
Proposition 3.1.
Let be such that Assumption B.1. Suppose that a -valued random function satisfies for all The identified set of is
It is easy to see that the identified set always contains a symmetric matrix, which is consistent with the consequence of Theorem 2.1. We also observe that cross-sectional demand distributions put no restriction on the off-diagonal entries except for an trivial restriction
4 Conclusion
In this paper, we addressed a question whether the symmetry of the average Slutsky matrix has empirical implications for cross-sectional demand distributions. We showed that the answer is negative: for any regular cross-sectional demand distributions, there exists a stochastic demand system that has a symmetric average Slutsky matrix. This result implies the untestability of individual Slutsky symmetry based on the average Slutsky matrix. Nonetheless, it does not rule out the possibility that individual Slutsky symmetry could have the empirical content through nonlinear relationship, which remains an important open question for future research.
Appendix A Preliminaries
Let be a path of probability measures on
Assumption A.1.
A.1.1
where is a bounded -domain.
A.1.2
is a -diffeomorphism satisfying
A.1.3
has density with respect to Lebesgue measure, and
A.1.4
There is such that
A.1.5
For each and is in for small
A.1.6
For each is in and
A.1.7
A.1.8
For and and hold.
Theorem A.1.
If satisfies Assumption A.1, then there exists a map such that it is continuously differentiable in both arguments on and where is a random variable drawn from
Proof.
For each define as Note that by (A.1.6). Consider the following Poisson equation with a Neumann boundary condition:
where is the outward unit normal vector on This PDE admits a unique solution such that by [nardi2015schauder]. Define a vector field as We first establish the regularity of
Lemma A.1.
The vector field
1.
is Lipschitz continuous uniformly over that is,
and
2.
is differentiable with a HΓΆlder continuous derivative uniformly over that is,
for some
Moreover, for and the map which is well-defined for small is continuous at
Lemma A.2.
There exists a continuous vector field such that it is Lipschitz in uniformly over and for
For simplicity, the extension is also denoted by We observe that solves the continuity equation
in the weak sense (see Chapter 4 of [santambrogio2015optimal]), i.e, for the map is absolutely continuous and it holds
Indeed, the fist condition follows because the map is an antiderivative of and the second holds because
where the second equality holds since the boundary integral vanishes due to the Neumann condition.
Consider the ODE
By Lemma A.2 and the Cauchy-Lipchitz theorem, there exists a unique solution that is in and its flow is a homeomorphism.
Let Then, satisfies the continuity equation in the weak sense by the standard argument. By the uniqueness of the solution of the continuity equation (Theorem 4.4 of [santambrogio2015optimal]), we have Hence, for a fixed random variable the random process has a continuously differentiable sample path and
We finally show that is continuously differentiable. Recall that is a homeomorphism between and Let Since is continuous by Lemma A.1, there exists a unique matrix-valued valued function that satisfies the linear ODE
(Theorem 3.9 of [teschl2012ordinary]). Let Letting
for we observe that
(Notice that for close enough to ) Since we have
By Gronwallβs inequality, we obtain
for some universal constant because is Lipschitz by Theorem 2.8 of [teschl2012ordinary] and is uniformly HΓΆlder by Lemma A.1. Hence, as which immediately implies is differentiable at with derivative
which is continuous in β
Next, we investigate whether the random function constructed in Theorem A.1 depends on the given path of probability measure smoothly. Let be an open interval. Let be a parametrized family of paths of probability measures on
Assumption A.2.
A.2.1
For each the path satisfies Assumption A.1.
A.2.2
is continuously differentiable.
A.2.3
Let and There exist functions and such that and hold.
Theorem A.2.
Proof.
By Theorem A.1, satisfies
where is the extension of by Lemma A.2. We first show that is smooth in the following sense.
Lemma A.3.
The vector field is differentiable in and and
1.
is continuous,
2.
is continuous,
3.
is continuous uniformly in and
4.
is continuous uniformly in
Consider the ODE
where
By Lemma A.3 and Theorem 3.12 of [teschl2012ordinary], this ODE admits a unique solution for each By Theorem 2.8 of [teschl2012ordinary], we have
for some constant Since the RHS converges to zero as by Lemma A.3, the map is continuous for each Finally, Lemma A.3 and Gronwallβs inequality show β
Appendix B Regularity Conditions on
Appendix C Omitted Proofs
Proof of lemma 3.1.
For simplicity, we consider the case of There exists a continuously differentiable function such that for by applying Theorem A.2 to the family Next, fix There exists a continuously differentiable function such that for by applying Theorem A.2 to the family Finally, fix There exists a continuously differentiable function such that for by applying Theorem A.2 to the family Let for and By construction, holds. The partial differentiability of in and follows by Theorem A.2. β
Proof of Lemma 3.2.
Recall that the average Slutsky matrix at is
The symmetry of this matrix is equivalent to
where the RHS is identified from the marginal distributions because
holds. Since
holds, the symmetry is equivalent to
which completes the proof. β
Proof of Lemma 3.3.
Let be a continuously differentiable function that has a compact support, is smooth in and
For given let
and
It is clear that in and on We also have
where the second equality holds by the integration by parts. Now, it is clear that satisfies the PDE (4). β
Proof of Lemma 3.4.
Since the second and third properties are obvious by construction, we shall prove the other two. The marginal compliance is shown using the continuity equation. Let be a test function. Then, the map is absolutely continuous and it holds
where the last equality holds by Lemma 3.3. For the partial differentiability, it suffices to show is differentiable, which is a consequence of Theorem A.2. Finally, the average symmetry condition (3) holds since
holds by Lemma 3.3 and the definition of β
Proof of Proposition 3.1.
Proof of Lemma A.1.
By the change-of-variable formula (see, for example, Lemma 2.62 of [sokolowski1992introduction]), solves
where and By (A.1.2), (A.1.6), and Theorem 2.1 of [kono2024well], is a unique solution satisfying and by his Theorem 3.1, an Schauder estimate
holds for some independent of 444Note that the coefficient is fixed in the original version of Theorem 3.1 of [kono2024well], but Theorem 6.30 of [gilbarg1977elliptic] ensures that depends on only through which is bounded by A.1.2. Using the equality we have
where is independent of which shows the uniform Lipschitzness of
Proof of Lemma A.2.
Let
which is finite by Lemma A.1. For define
for Since is continuous, so is by Bergeβs maximum theorem. Also, McShaneβs extension theorem implies that is Lipschitz uniformly over β
Proof of Lemma A.3.
The differentiability of in is obvious. The first and third bullet points are shown in the same way as Lemma A.1. To show the second and fourth points, observe that the following PDE holds:
The standard Schauder estimate shows the differentiability of by Assumption A.2.3. Since the derivative is continuous in and by Assumption A.2.3, the second and fourth bullet points are confirmed. β