### v2

# Untestability of Average Slutsky Symmetry

Abstract
Slutsky symmetry and negative semidefiniteness are necessary and sufficient conditions for the rationality of demand functions. While the empirical implications of Slutsky negative semidefiniteness in repeated cross-sectional demand data are well understood, the empirical content of Slutsky symmetry remains largely unexplored. This paper takes an important first step toward addressing this gap. We show that the average Slutsky matrix is not identified and that its identified set always contains a symmetric matrix, implying that the symmetry of the average Slutsky matrix is untestable and that individual Slutsky symmetry cannot be tested through the average. Nevertheless, we demonstrate that, by imposing bounds on the income elasticity of demand, Slutsky symmetry implies a set of functional inequality constraints that are testable.
Keywords— Slutsky symmetry, rationality, cross-sectional demand, continuity equation
1 Introduction
Rationality is a central concept in economic theory, serving as a fundamental assumption in the analysis of consumer behavior. It assumes that consumers make decisions that are consistent with utility maximization, a principle that supports many economic models and empirical research. Testing this assumption is crucial for validating theoretical models and understanding consumer decision-making.
When individual demand functions are available, the Hurwicz-Uzawa theorem ([8]) provides a complete characterization of rationality. To state the theorem, let be a -dimensional price vector and be an income, both of which are relative to the price of the numeraire. Let be the -dimensional vector of quantities demanded when the price is and the income is We say that a demand function is rational if it maximizes a utility function subject to the budget constraint. According to the Hurwicz-Uzawa theorem, a demand function is rational if and only if its Slutsky matrix, which is defined as a matrix
is both symmetric and negative semidefinite.
In many real-world applications, individual demand functions are not observed and only cross-sectional data is available. To investigate the population rationality, researchers have explored the rationalizability of the average demand function conditional on observable characteristics, such as price and income, treating it as if it is generated by an individual representing the population. See, for example, [13], [14], [5], and [7].
However, this approach involves some disadvantages. For instance, as Theorem 1 of [14] demonstrates, the rationality of the average demand function is irrelevant to the rationality of individual demand functions: it is possible that all individuals in a population are rational while the average demand is not, and vice versa. Moreover, it is often reported that cross-sectional mean regressions fail to explain the variation of demand adequately because of the unobserved preference heterogeneity. ([7], [6]). These facts indicate the need to investigate all the information about the heterogeneity contained in the data.
Compared with testing the rationality of an individual demand function, testing whether a cross-sectional demand data is consistent with a rational population is a harder problem because individual Slutsky matrices are not observed. This difficulty raises a fundamental question: can we infer the rationality of a population from cross-sectional demand data? More specifically, is it possible to construct a statistical test with nontrivial power to detect consumer irrationality only from demand distributions?
Recent research has made progress in addressing this question. [6] provide a necessary and sufficient condition for cross-sectional demand distributions to be rationalizable when only two goods are present. For cases involving more than two goods, [2] provide the empirical content of the negative semidefiniteness of the Slutsky matrix. Specifically, they show that the quadratic form of the average Slutsky matrix conditional on price and income, where is the random demand function representing the population, is identified. If an estimate of the quadratic form is significantly positive, then it means that the observed data is likely to be inconsistent with a population having an negative semidefinite Slutsky matrix—let alone rationality. These studies demonstrate the feasibility of testing rationality by means of the negative semidefiniteness of the Slutsky matrix.
However, these methods neglect the symmetry condition, leading to statistical tests that are overly conservative and fail to fully leverage the implications of rationality. The symmetry condition can be more critical to the rationality than the negative semidefiniteness, given that the former is highly sensitive to small perturbations while the latter is not. If Slutsky symmetry is testable, one should be able to construct a much more powerful test for rationality by testing both conditions.
Despite its theoretical importance, the empirical content of the Slutsky symmetry condition remains poorly understood. This paper provides a first step toward elucidating the implications of individual Slutsky symmetry for cross-sectional demand distributions. The main result of this paper claims that the average Slutsky matrix is not identified, even though its quadratic form is identified, as shown in [2]. Moreover, we show that the identified set of the average Slutsky matrix always contains a symmetric matrix.
An immediate consequence of our results is the fundamental untestability of individual Slutsky symmetry via the average Slutsky matrix. This fact indicates that the Slutsky symmetric condition is totally distinct to the negative semidefiniteness condition, which is testable based on the quadratic form of the average Slutsky matrix ([2]).
The untestability of the average Slutsky symmetry does not immediately rule out the possibility of testing Slutsky symmetry from cross-sectional distributional data. After the first draft of this paper was made public, [4] derive the empirical content of individual Slutsky symmetry without relying on the average Slutsky matrix. They show that individual Slutsky symmetry implies a nonparametric conditional quantile restriction, which is testable.
Although the average Slutsky symmetry is untestable in general, it can become testable under additional assumptions. In particular, we consider imposing bounds on the income elasticity of demand. By restricting the patterns of demand substitution in this way, Slutsky symmetry leads to functional inequality conditions that are, in principle, testable. Constructing a formal statistical test based on these conditions is beyond the scope of this paper, but existing methods for testing functional inequalities could be applied.
A key theoretical contribution of this paper is the development of a constructive method to generate a random demand function that satisfies prescribed marginal distributions while remaining consistent with the Slutsky conditions. In general, both the levels and the derivatives of individual demand functions vary across consumers, reflecting heterogeneous preferences. The construction in this paper shows, however, that even if one fixes a deterministic rule describing how demand changes with prices and income, it is still possible to produce random demand distributions that are consistent with observed marginals and with rational behavior. The key insight is that once the local response of demand to prices and income is fixed, the distribution of demand at all other price-income pairs can be obtained from the distribution at a single reference point by following the trajectories implied by this response rule. Mathematically, this propagation can be interpreted as the solution to a system of differential equations that transports the reference distribution along price-income changes.
Related literature. [8] investigate the integrability of individual demand functions and give a necessary and sufficient condition for rationality based on the Slutsky matrix. [14] considers a population that is heterogeneous in preference and provides conditions for the average demand function to be rational. [5] and [7] among others consider the estimation and inference on the average demand function under the rationality shape restriction.
For cross-sectional demand distributions, [6] give a necessary and sufficient condition for observed datasets to be consistent with a rational demand system for the cases when there are only two goods. When more than two goods are present, the characterization of rationalizability is largely open. [2] provide a necessary condition that the data needs to satisfy for it to be rationalizable focusing on the negative semidefiniteness of the Slutsky matrix and propose a statistical testing for rationality. [16] construct a testing procedure of rationality based on higher order moments of demand distributions. They also observe that the average Slutsky matrix is not identified from those moments. More recently, [4] derive the empirical content of individual Slutsky symmetry without relying on the average Slutsky matrix.
This paper is also related to the literature on random utility models. [18] and [17] show that the axiom of revealed stochastic preference characterizes rationalizability of stochastic choice functions defined on a finite number of choice sets. In a similar setup, [9] constructed a statistical test for the axiom. In the context of discrete choice, [1] recently gives the complete characterization of the rationalizability of demand distributions.
2 Setup and Results
2.1 Setup
For we consider an economy with goods. Relative to the first good, their prices are encoded into a price vector where Let be the relative income. For given price and income a consumer demands Notice that we assume the homogeneity of demand functions at this point. We also assume Warlas’ law, i.e., the demand for the numeraire is and consequently, is assumed.
We assume
for and Let We restrict ourselves to demand functions that are in the space
In the real world, consumer’s preference heterogeneity is present. In this sense, consumer’s demand is stochastic from the perspective of an econometrician. Let be a random individual demand function drawn from a probability distribution on The econometrician is assumed to observe cross-sectional demand distributions, that is, (s)he observes the distribution of for each but no joint distribution of demands across different price-income levels is available. Regarding as a stochastic process indexed by we often call the (one-dimensional) marginal distribution of at Note that although is a population object, we assume that it is available since we are interested in identification. Assume that the interior of the support of denoted by is not empty.
Example 1.
For the three-good case (), consider a random Cobb-Douglas demand where is a random vector such that and Then is a -valued random element. The demand distribution conditional on is a distribution supported on a subset of the triangle generated by and
Recall that an individual demand function is said rational if it is induced by utility maximization. The Hurwicz-Uzawa theorem states that is rational if and only if its Slutsky matrix
is symmetric and negative semidefinite for each
2.2 Main Result
As we discussed in the previous section, [2] propose a method to test individual rationality based on the negative semidefiniteness. Their testing procedure is roughly as follows. In their Theorem 1, they show that the quadratic form of the average Slutsky matrix is identified, that is, is identified for each and If the population consists of rational individuals, then is negative semidefinite almost surely, and hence, should hold. One can reject the null hypothesis that the population is rational if an estimate of is significantly positive for some and 222To be more precise, [2] show a stronger result that the quadratic form of the average Slutsky matrix conditional on the value of Thus, their test is more powerful than what is described here.
Although this method has nontrivial power, it is likely to be overly conservative since it completely neglects the other critical component of rationality, Slutsky symmetry. We shall investigate its testability based on the average Slutsky matrix. This question is important since the symmetry property is not robust to small perturbations while the negative semidefiniteness is. If Slutsky symmetry is testable, one should be able to construct a much more powerful test for rationality by testing both conditions.
If individual Slutsky matrix is symmetric (almost surely), so is its average. In what follows, we address the question of what we can say about the average Slutsky matrix from cross-sectional demand distributions
First, we observe that the average Slutsky matrix is not identified in an explicit way because of the second term of (1). The expectation of the second term is written as
but it is unclear how to identify because it involves the joint distribution of demand at different income levels, and it is indeed not identifiable as Theorem 2.1 implies below.
The nonidentifiability of the average Slutsky matrix does not immediately imply that we cannot say anything about its symmetry. The individual symmetry could have some empirical implications for the observable demand distributions. Unfortunately, however, there is nothing we can say about the symmetry of the average Slutsky matrix from the cross-sectional demand data. The following theorem formalizes the argument so far. The proof is given in Appendix.
Theorem 2.1.
Let be such that Assumption B.1. The identified set of the function that maps price-income pairs to the average Slutsky matrix at is given by
where
is identified. In particular, the average Slutsky matrix is not identified at any and there necessarily exists a -valued random demand function such that and is symmetric for all
Assumption B.1 imposes regularity on the family of demand distributions. Specifically, it requires that has a smooth density, and that both the density and support vary smoothly with
Remark 2.1.
Proposition 2 of [16] asserts that the average Slutsky matrix is not “automatically” identified. They explain that the quantity (2) is not identified in the same way that the other term is identified through the equation Their argument is not complete, as it merely rules out a particular strategy for identifying the average Slutsky matrix without addressing the possibility of alternative identification approaches. Consequently, [16] do not establish whether there exists an observationally equivalent demand system whose average Slutsky matrix is symmetric.
2.3 Practical Implications
Testability of Slutsky symmetry. Theorem 2.1 has several useful implications. First, as stated in the theorem, the average Slutsky matrix is not identified, and the identified set is unbounded in the sense that can be arbitrarily large by is small. Moreover, Theorem 2.1 implies that no matter what cross-sectional demand distributions satisfying the regularity condition we observe, there always exists a stochastic demand system such that it is observationally equivalent to the true demand system and its average Slutsky matrix is symmetric. Consequently, there is no way to infer whether is symmetric or not from
This negative result suggests that researchers collect additional demand data in order to identify the average Slutsky matrix and test Slutsky symmetry. For example, let us assume that the joint distribution of is available for all rather than the marginal distribution This setup corresponds to the situation where analysts can observe each individual’s choice twice. In this setup, the average Slutsky matrix is identified because the RHS of (2) is identified, and hence, it is possible to test the Slutsky symmetry by testing whether is symmetric or not.
It is also worth mentioning that even though Theorem 2.1 does show that Slutsky symmetry is not testable based on the average Slutsky matrix, it does not immediately rule out the possibility of testing Slutsky symmetry from cross-sectional distributional data. Indeed, individual symmetry could have observable implications through nonlinear statistics rather than the average, but we leave this for future work.
Adding income elasticity bounds. Now, we shall see that the average Slutsky symmetry can be testable by imposing bounds on the income elasticity of demand. Let
be the income elasticity of demand for th good for demand system Then, we have
Let be the income elasticity of demand It is easy to check
It is often reasonable to put bounds on the income elasticity of demand Assume that there are functions such that for all and Then we have
This inequality leads to a bound on the difference for it must be lie in the interval
Since is identified and estimable, the average Slutsky symmetry— for all and —can be tested by checking whether the null hypothesis
While constructing a statistical test for this hypothesis is beyond the scope of this paper, the literature on testing functional inequalities, such as [11], [12], and [15], will work.
3 Construction of Stochastic Demand Systems
Let be an element of the identified set of the average Slutsky matrix. To prove Theorem 2.1, it is sufficient to construct a -valued random demand function such that
for The goal of this section is to describe the construction of such a random demand function.
In overview, the construction proceeds in two steps. First, we construct a preliminary random demand function such that (M) holds but not necessarily (S). In the second step, we modify to obtain another random demand function that satisfies (S) as well as (M).
Step 1. We begin with constructing a preliminary random demand function satisfying (M). Let
Lemma 3.1.
Remark 3.1.
In the random demand function constructed in Lemma 3.1, consumers’ preference heterogeneity, or “type,” is encoded in By the third property of consumer’s type is understood as the demand at Observe that the demand system constructed in Lemma 3.1 is degenerated in the sense that consumers are completely characterized by the demand at More specifically, if an individual demands at (s)he demands at for sure.
Before moving on to the next step, we observe that the random demand function is characterized by the solution to an ordinary differential equation (ODE). Fix Let be the income derivative of the demand function of consumer
where Notice that this is well-defined since is homeomorphic. The family of vector fields, combined with an initial condition, pins down the demand function, as the Cauchy-Lipschitz theorem implies that is the unique solution to the ODE
under the global Lipschitz condition on
Step 2. Although in Lemma 3.1 satisfies the marginal compliance (M), it does not satisfy the average Slutsky symmetry condition in general. Our strategy is to construct another random demand function satisfying both conditions by modifying To do so, we modify the ODE (3) so that its solution respects (S). More precisely, we rectify by adding an auxiliary vector field constructed in the following lemma.
Lemma 3.2.
For let be a continuous function on Assume that is satisfied for all There exists a vector field that is Lipschitz uniformly in and satisfies
For set
which is a continuous function on under Assumption B.1. Then we have if since and
where the first equality holds by the law of motion of the ODE (3), and the second equality holds by the fact that satisfies (M). Take a vector field from Lemma 3.2 for this and let Consider a modified ODE
By the Cauchy-Lipchitz theorem, this ODE admits a unique solution for each Define
The following lemma shows that this flow induces a random demand function that satisfies the desired requirements.
Lemma 3.3.
4 Conclusion
In this paper, we have shown that the average Slutsky symmetry is not testable using cross-sectional demand data. To establish this, we explicitly derived the identified set of the average Slutsky matrix and demonstrated that it always contains a symmetric matrix. This finding implies that individual Slutsky symmetry cannot be tested through the average Slutsky matrix, although it does not rule out the possibility of testing the symmetry hypothesis without relying on the average. See [4] for recent developments. Furthermore, by imposing bounds on the income elasticity of demand, we showed that average Slutsky symmetry leads to a set of functional inequality constraints that are, in principle, testable. A promising direction for future research is to develop a statistical test for Slutsky symmetry based on these inequalities.
Appendix A Preliminaries
Let be a path of probability measures on
Assumption A.1.
A.1.1
where is a bounded -domain.
A.1.2
is a -diffeomorphism satisfying
A.1.3
has density with respect to Lebesgue measure, and
A.1.4
There is such that
A.1.5
For each and is in for small
A.1.6
For each is in and
A.1.7
A.1.8
For and and hold.
Theorem A.1.
If satisfies Assumption A.1, then there exists a map such that it is continuously differentiable in both arguments on and where is a random variable drawn from
Proof.
For each define as Note that by (A.1.6). Consider the following Poisson equation with a Neumann boundary condition:
where is the outward unit normal vector on This PDE admits a unique solution such that by [19]. Define a vector field as We first establish the regularity of
Lemma A.1.
The vector field
1.
is Lipschitz continuous uniformly over that is,
and
2.
is differentiable with a Hölder continuous derivative uniformly over that is,
for some
Moreover, for and the map which is well-defined for small is continuous at
Lemma A.2.
There exists a continuous vector field such that it is Lipschitz in uniformly over and for
For simplicity, the extension is also denoted by We observe that solves the continuity equation
in the weak sense (see Chapter 4 of [20]), i.e, for the map is absolutely continuous and it holds
Indeed, the fist condition follows because the map is an antiderivative of and the second holds because
where the second equality holds since the boundary integral vanishes due to the Neumann condition.
Consider the ODE
By Lemma A.2 and the Cauchy-Lipchitz theorem, there exists a unique solution that is in and its flow is a homeomorphism.
Let Then, satisfies the continuity equation in the weak sense by the standard argument. By the uniqueness of the solution of the continuity equation (Theorem 4.4 of [20]), we have Hence, for a fixed random variable the random process has a continuously differentiable sample path and
We finally show that is continuously differentiable. Recall that is a homeomorphism between and Let Since is continuous by Lemma A.1, there exists a unique matrix-valued valued function that satisfies the linear ODE
(Theorem 3.9 of [22]). Let Letting
for we observe that
(Notice that for close enough to ) Since we have
By Gronwall’s inequality, we obtain
for some universal constant because is Lipschitz by Theorem 2.8 of [22] and is uniformly Hölder by Lemma A.1. Hence, as which immediately implies is differentiable at with derivative
which is continuous in ∎
Next, we investigate whether the random function constructed in Theorem A.1 depends on the given path of probability measure smoothly. Let be an open interval. Let be a parametrized family of paths of probability measures on
Assumption A.2.
A.2.1
For each the path satisfies Assumption A.1.
A.2.2
is continuously differentiable.
A.2.3
Let and There exist functions and such that
hold.
Theorem A.2.
Appendix B Regularity Conditions on
Appendix C Omitted Proofs
Proof of Lemma 3.1.
For simplicity, we consider the case of There exists a continuously differentiable function such that for by applying Theorem A.2 to the family Next, fix There exists a continuously differentiable function such that for by applying Theorem A.2 to the family Finally, fix There exists a continuously differentiable function such that for by applying Theorem A.2 to the family Let for and By construction, holds. The partial differentiability of in and follows by Theorem A.2. ∎
Proof of Lemma 3.2.
Let be an open ball such that in There exists a smooth nonnegative function such that and For each define a function as Then is and satisfies Moreover, the derivative is Lipschitz uniformly in due to the regularity of If we define then holds. Let where
We shall see that satisfies all the conditions of Lemma 3.2. By the property of and continuity of is Lipschitz uniformly in The third condition of (4) is satisfied because for
where the second equality holds by the integration by parts, and similarly, for
The second condition holds since vanishes on the boundary. The first condition holds since
Proof of Lemma 3.3.
For the partial differentiability, it suffices to show is differentiable, which is a consequence of Theorem A.2. The marginal compliance (M) is shown using the continuity equation. Let be a test function. Then, the map is absolutely continuous and it holds
where the last equality holds by Lemma 3.2. By Lemma 3.2 and the definition of we have
which implies that condition (S) holds. ∎
Proof of Theorem 2.1.
It is easy to see the function is an element of Conversely, Lemma 3.3 shows that for a given there exists a -valued random demand function such that and for all For the second part of the theorem, if we define
then holds, and is symmetric for all ∎
Proof of Lemma A.1.
By the change-of-variable formula (see, for example, Lemma 2.62 of [21]), solves
where and By (A.1.2), (A.1.6), and Theorem 2.1 of [10], is the unique solution satisfying and by his Theorem 3.1, an Schauder estimate
holds for some independent of 333Note that the coefficient is fixed in the original version of Theorem 3.1 of [10], but Theorem 6.30 of [3] ensures that depends on only through which is bounded by A.1.2. Using the equality we have
where is independent of which shows the uniform Lipschitzness of
Proof of Lemma A.2.
Let
which is finite by Lemma A.1. For define
for Since is continuous, so is by Berge’s maximum theorem. Also, McShane’s extension theorem implies that is Lipschitz uniformly over ∎
Proof of Lemma A.3.
The differentiability of in is obvious. The first and third bullet points are shown in the same way as Lemma A.1. To show the second and fourth points, observe that the following PDE holds:
The standard Schauder estimate shows the differentiability of by Assumption A.2.3. Since the derivative is continuous in and by Assumption A.2.3, the second and fourth bullet points are confirmed. ∎